OpenStudy (anonymous):

The temperature of an oven went from 360° to 280°. What was the percent decrease in temperature to the nearest percent? A. 22% B. 29% C. 80% D. 130%

OpenStudy (anonymous):

@ROBLOX

OpenStudy (anonymous):

initial-final/ initial 360-280/ 360 22%

OpenStudy (welshfella):

first worl out the difference 360 - 280 = ?

OpenStudy (mathstudent55):

\(percent~change = \dfrac{new~value - old~value}{old~value} \times 100\)

OpenStudy (mathstudent55):

When the percent change is negative, it is a percent decrease. When it is positive, it is a percent increase.

OpenStudy (anonymous):

it's c i know

OpenStudy (anonymous):

it is b

OpenStudy (anonymous):

360-280 is 80 then 80/360 is .22 then .22 x 100 is 22%

OpenStudy (anonymous):

no it's c i'm sure trust me

OpenStudy (anonymous):

do this equation \[\frac{ 280-360 }{280} \times 100\]

OpenStudy (anonymous):

then round to the nearest percent

OpenStudy (anonymous):

it would be divided by 360 but ok :)

OpenStudy (anonymous):

well divideding would be 1.285714286

OpenStudy (anonymous):

right

OpenStudy (welshfella):